# Carbon-Aware Optimal Power Flow

### v2

@@ I Introduction
In practice, there are two main types of methods currently used for demand-side carbon accounting [3]: 1) the location-based method, which employs grid average emission factors (AEFs) across long time horizons and large areas [6, 3] to account for electricity users’ carbon footprints, and 2) the market-based method, which derives carbon emission factors purely based on clean power market instruments [7], such as Renewable Energy Certificates (RECs) [8] and Power Purchase Agreement (PPA) [9]. See [4] for a detailed introduction. This paper focuses on the location-based method, and the existing AEF-based carbon accounting schemes suffer from two critical limitations: 1) lack of temporal and spatial granularity, and 2) disregard of actual electricity delivery through physical power networks [4]. Since the generation fuel mix in power systems constantly changes over time, the grid carbon intensities of electricity are dynamic and time-varying with significant daily and seasonal patterns [10]. In addition, the grid carbon intensities vary geographically, as generators of different fuel types are distributed across various locations. Moreover, power grids feature specific network topologies and circuits through which power flows, physically connecting end-users with generators and impacting carbon emissions. Therefore, carbon accounting schemes require sufficient granularity and alignment with physical power grids to 1) reflect the temporal variations and spatial diversity in grid emissions; 2) provide precise and accurate carbon accounting results for end-users; and 3) effectively inform and incentivize grid decarbonization decisions. See [4, 6, 11] for more discussions and justifications.
@@
To tackle these issues, the concept of carbon (emission) flow is introduced in [12], where carbon emissions are treated as virtual network flows embodied in energy flows, transmitted from producers to consumers. References [13, 14] establish the mathematical models of carbon flow in electric power networks. The carbon flow method defines nodal and branch carbon intensities for the grid, providing a temporally and spatially granular depiction of the grid’s carbon footprints. In this way, the carbon flow method represents a prominent tool for demand-side carbon accounting, which aligns carbon footprint calculations with the physical grids and power flows. See [13] and Section II for more details.
@@
This paper is dedicated to advancing fundamental theories and methodologies for decarbonization decision-making in power systems. In this regard, the Optimal Power Flow (OPF) method [15, 16] stands as a foundational mathematical tool for optimizing power system decisions. It has been studied extensively in the literature and widely applied in grid planning, operations, control, and electricity markets [17, 18, 19]. Existing OPF schemes typically seek optimal power grid decisions to minimize a specific economic cost objective, while satisfying the power flow equations, network constraints, and operational limits of devices. However, the essential goals of grid decarbonization and carbon footprint management are inadequately considered. Since power system decisions, e.g., the siting and sizing of renewable generators or power dispatch of various generation sources, directly impact the grid’s carbon footprints, it becomes necessary to explicitly integrate carbon footprint management into grid decision-making for achieving desired decarbonization performances and outcomes.
@@ I-A Related Work and Key Issues
    (Power Flow Directions): The carbon flow model [13] needs to pre-determine the power flow directions for all branches to identify the power inflows for each node. However, the directions of branch power flows are typically unknown prior to solving optimal decision models. Most existing works that employ the carbon flow model in grid decision-making either overlook the issue of unknown power flow directions or assume they are predetermined by alternating between grid optimization and carbon flow calculation through iterations [29]. Reference [25] introduces binary indicator variables to handle the unknown power flow directions in the carbon flow model. It results in a mixed-integer nonconvex quadratically constrained optimization problem with both integer variables and nonconvex constraints. Thus, a tailored heuristic penalty-based iterative algorithm is designed to solve the optimization problem via iterations.
@@
    (Carbon Footprint Model for Energy Storage): Due to the switching between charging and discharging, energy storage (ES) systems can shift the grid’s carbon footprints across time, resulting in time-coupled carbon flows. The carbon accounting for ES systems is an emerging topic that is attracting considerable recent attention from the industry [31, 32]. References [26, 25, 33] propose different carbon footprint models for ES systems, while they neglect the carbon emission leakage associated with ES energy loss (see Remark 4), and the carbon accounting mechanisms for ES system owners remain unclear.
@@ II-B Carbon Emission Flow Model for Lossy Power Networks
Equation (4) is referred to as the carbon flow model or carbon flow equation, and it can be equivalently reformulated in the matrix form (5). The detailed derivation of (5) is provided in Appendix A.
$(\bm{P}_{\mathrm{N}}-\bm{P}_{\mathrm{B}})\bm{w}_{\mathrm{N}}=\bm{r}_{\mathrm{G}}\Longrightarrow\bm{w}_{\mathrm{N}}=(\bm{P}_{\mathrm{N}}-\bm{P}_{\mathrm{B}})^{-1}\bm{r}_{\mathrm{G}},$ (5)
Here, $\bm{w}_{\mathrm{N}}\!:=\!(w_{i})_{i\in\mathcal{N}}\!\in\!\mathbb{R}^{N}$ and $\bm{r}_{\mathrm{G}}\!:=\!(\sum_{g\in\mathcal{G}_{i}}\!\!w_{i,g}^{\mathrm{G}}P_{i,g}^{\mathrm{G}})_{i\in\mathcal{N}}\!\in\!\mathbb{R}^{N}$ denote the column vectors that collect the nodal carbon intensities and nodal generation carbon flows, respectively. $\bm{P}_{\mathrm{N}}\!:=\!\text{diag}(P_{i}^{\mathrm{in}})\!\in\!\mathbb{R}^{N\times N}$ is the diagonal matrix whose $i$-th diagonal entry is the nodal active power inflow $P_{i}^{\mathrm{in}}$ to node $i$. $\bm{P}_{\mathrm{B}}\!\in\!\mathbb{R}^{N\times N}$ is the branch power inflow matrix that is built by letting $\bm{P}_{\mathrm{B}}[i,k]\!=\!P_{ki}^{i}$ and $\bm{P}_{\mathrm{B}}[k,i]\!=\!0$ if node $k$ sends power flow $P_{ki}^{i}$ to node $i$. Note that $P_{ki}^{i}$ and $P_{ki}^{k}$ are different due to the line power loss. See [13] for more explanations.
@@ II-C Feasibility and Solution Uniqueness of Carbon Flow Model
According to [36, Theorem], the matrix $\bm{P}_{\mathrm{C}}$ is nonsingular and thus invertible. Since $\bm{P}_{\mathrm{C}}$ is diagonally dominant and all its off-diagonal elements $\{\bm{P}_{\mathrm{C}}[i,j]\}_{i\neq j}$ are non-positive, $\bm{P}_{\mathrm{C}}$ is an M-matrix [37] according to [36, Corollary 4]. ∎
@@
The condition in Theorem 1 can be interpreted as that for any node $i$ with no generation power injection, one can find a power flow path $i\leftarrow i_{1}\leftarrow i_{2}\leftarrow\cdots\leftarrow i_{r}\leftarrow j$ in the power network that traces upstream to a node $j$ that has generation power injection. This condition generally holds for practical connected power networks. This condition also implies that every node has non-zero power flux, i.e., $\bm{P}_{\mathrm{C}}[i,i]>0$ for all $i\in[N]$. However, non-zero nodal power flux can not guarantee that $\bm{P}_{\mathrm{C}}$ is invertible. For example, consider a simple 3-node circular network case shown in Figure 3. In this case, every node has 1 unit of power flowing through it, but the matrix $\bm{P}_{\mathrm{C}}$ is singular. And this case violates the condition in Theorem 1. In addition, $\bm{P}_{\mathrm{C}}$ possesses all the properties of being an M-matrix. For example, all the elements of $\bm{P}_{\mathrm{C}}^{-1}$ are non-negative and every eigenvalue of $\bm{P}_{\mathrm{C}}$ has positive real part [37, Theorem 1.1].
@@ III-A1 Objective Function
where (8a) denotes the total generation cost in a quadratic form with the parameters $c_{i,g}^{2},c_{i,g}^{2},c_{i,g}^{0}$, and (8b) is the penalty on generation-side emissions with the cost coefficient $c^{\mathrm{emi}}$.
@@ III-A2 Power Flow Equations and Constraints
The power flow constraints (7c) generally involve the line thermal constraints (10a) and voltage limits (10b).
@@
where $\bar{S}_{ij}$ is the line apparent power capacity, and $\bar{V}_{i},\underline{V}_{i}$ are the upper and lower limits of voltage magnitude at node $i$.
@@ III-A3 Carbon Flow Equations and Constraints
Essentially, the C-OPF model (7) is a carbon-aware generalization of the OPF model, and it reduces to an OPF model if the carbon-related objective function $f_{\mathrm{carbon}}$, carbon flow equations (7d) and constraints (7e) are removed or inactive. It also implies that existing OPF techniques, such as linearization, convexification, decomposition, stochastic modeling, etc., can still be applied to the power flow components in C-OPF (7). Moreover, the C-OPF model (7) can extend to the multi-time period setting with time-coupling constraints. As an illustrative example, a complete C-OPF-based economic dispatch model (27) is established in Section V-A.
@@ IV Carbon Footprint Model for Energy Storage
Energy storage (ES) systems play a pivotal role in decarbonizing power grids, as their operations can be optimized to curtail overall system emissions, e.g., charging when the grid is clean and discharging when the grid is under high emissions. Consider an ES system connected to node $i\!\in\!\mathcal{N}$. For time $t\in\mathcal{T}:=\{1,2,\cdots,T\}$ with the time interval $\delta_{t}$, the dynamical ES power model [42] is formulated as (20):
@@ V-C1 Nodal Carbon Intensity
Figure 6 illustrates the nodal carbon intensities for all load nodes under the C-OPF-based and OPF-based ED schemes, respectively. It shows that the C-OPF-based ED model can generate effective power dispatch schemes that keep the nodal carbon intensities of load nodes below the cap of $\bar{w}_{i,t}\!=\!1.2$ lbs/kWh. In contrast, the OPF-based ED scheme frequently exceeds the nodal carbon intensity cap. Moreover, the grid’s nodal carbon intensities exhibit significant temporal variation and spatial diversity. From Figure 6, it is observed that the nodal carbon intensities at 13:00 are generally lower than those at 21:00, due to the higher penetration of renewable generation at 13:00, as illustrated in Figure 8. Figure 7 visualizes the grid’s nodal carbon intensities at 13:00, with darker blue colors indicating higher carbon intensity at each node. It demonstrates that the nodal carbon intensities calculated using the carbon flow method can reflect the proximity to different fuel types of generation and align with the physical power flow. In comparison, the grid average emission factor at 13:00 is calculated to be 0.5 lbs/kWh, which only gives an indication of the overall grid average emission profile and falls short of providing detailed insight into local emissions at different locations.
@@ V-F Computational Efficiency
Due to the dual power flow reformulation and the addition of carbon flow equations and constraints, the C-OPF model generally has a larger problem size and requires more solution time than the traditional OPF model. In practice, several methods can improve the solution efficiency of the C-OPF model. For instance, by inspecting all branches and identifying those whose power flow directions can be predetermined based on the network configuration, the dual power flow reformulation can be avoided for these branches. Additionally, the solutions of the OPF model can be used to warm-start the C-OPF model. Furthermore, a key future direction is to develop efficient linearization and convexification approaches for the nonconvex carbon flow equations to fundamentally enhance the solution efficiency of the C-OPF model.

### v4

@@ I Introduction
As introduced in [4], carbon accounting frameworks can be categorized into two primary types: attributional and consequential. Attributional carbon accounting aims to allocate direct generation-side emissions to end-users for assigning emissions responsibility, namely the Scope 2 emissions described above. In contrast, consequential carbon accounting seeks to evaluate the change or impact on grid emissions resulting from specific decisions or projects, compared to a counterfactual baseline emission scenario, which employs methods such as marginal emissions [6] and avoided emissions. These two carbon accounting frameworks are distinct and serve different purposes [4]. In terms of attributional carbon accounting, there are two main currently-used methods [3]: 1) the location-based method, which adopts grid average emission factors (AEFs) across long time horizons and large areas [7, 3] to account for electricity users’ carbon footprints, and 2) the market-based method, which derives carbon emission factors purely based on clean power market instruments [8], such as Renewable Energy Certificates (RECs) [9] and Power Purchase Agreement (PPA) [10]. See [4] for a comprehensive overview of carbon accounting methods for power grids.
@@
This paper focuses on the location-based method for attributional demand-side carbon accounting, while the existing AEF-based schemes suffer from two critical limitations: 1) lack of temporal and spatial granularity, and 2) disregard of actual electricity delivery through physical power networks [4]. Since the generation fuel mix in power systems constantly changes over time, the grid carbon intensities of electricity are dynamic and time-varying with significant daily and seasonal patterns [11]. In addition, the grid carbon intensities vary geographically, as generators of different fuel types are distributed across various locations. Moreover, power grids feature specific network topologies and circuits through which power flows, physically connecting end-users with generators and impacting carbon emissions. Therefore, carbon accounting schemes require sufficient granularity and alignment with physical power grids to 1) reflect the temporal variations and spatial diversity in grid emissions, 2) provide precise and accurate carbon accounting results for end-users, and 3) effectively inform and incentivize grid decarbonization decisions. See [4, 7, 12] for more discussions and justifications.
@@
To tackle these issues, the concept of carbon (emission) flow is introduced in [13], where carbon emissions are treated as virtual network flows embodied in energy flows, transmitted from producers to consumers. References [14, 15] establish the mathematical models of carbon flow in electric power networks. The carbon flow method defines nodal and branch carbon intensities for the grid, providing a temporally and spatially granular depiction of the grid’s carbon footprints. In this way, the carbon flow method represents a prominent tool for demand-side carbon accounting, which aligns carbon footprint calculations with the physical grids and power flows. See [14] and Section II for more details. Recent work [16] presents a tree search algorithm to trace the contribution of each generator to individual lines and loads, thereby estimating nodal carbon emissions. This approach builds on the concept of electricity flow tracing [17] and yields nodal carbon emission estimates comparable to those produced by the carbon flow method, with the potential distinction being in how power losses are handled.
@@
Unlike the studies above that focus on accounting for and estimating the grid carbon footprints, this paper aims to advance foundational methodologies for grid decarbonization decision-making and enable optimal carbon-electricity joint management in power grids. In this regard, the Optimal Power Flow (OPF) method [18, 19] stands as a foundational mathematical tool for optimizing power system decisions. It has been studied extensively in the literature and widely applied in grid planning, operations, control, and electricity markets [20, 21, 22]. Existing OPF schemes typically seek optimal power grid decisions to minimize a specific economic cost objective, while satisfying the power flow equations, network constraints, and operational limits of devices. However, the essential goals of grid decarbonization and carbon footprint management are inadequately considered. Since power system decisions, e.g., the siting and sizing of renewable generators or power dispatch of various generation sources, directly impact the grid’s carbon footprints, it becomes necessary to explicitly integrate carbon footprint management into grid decision-making for achieving desired decarbonization performances and outcomes.
@@ I-A Related Work and Key Issues
    (Power Flow Directions): The carbon flow model [14] needs to pre-determine the power flow directions for all branches to identify the power inflows for each node. However, the directions of branch power flows are typically unknown prior to solving optimal decision models. Most existing works that employ the carbon flow model in grid decision-making either overlook the issue of unknown power flow directions or assume they are predetermined by alternating between grid optimization and carbon flow calculation through iterations [32]. Reference [28] introduces binary indicator variables to handle the unknown power flow directions in the carbon flow model. It results in a mixed-integer nonconvex quadratically constrained optimization problem, which cannot be readily solved using off-the-shelf optimizers due to involving both integer variables and nonconvex constraints. Thus, a tailored heuristic penalty-based iterative algorithm is designed to solve the optimization problem.
@@
    (Carbon Footprint Models for Energy Storage): By switching between charging and discharging, energy storage (ES) systems can shift load demand and transfer renewable energy across time, offering substantial potential to reduce power system emissions. Hence, developing carbon footprint models for ES systems is crucial, since it lays the quantitative foundation for making optimal ES planning and operation decisions, such as determining installation location and capacity, and managing charging and discharging. In addition, granular and accurate carbon accounting for ES systems enhances information transparency, supporting the development of low-carbon regulatory policies and market mechanisms for ES. It also incentivizes ES stakeholders to adopt low-carbon practices to reduce their own carbon footprint and overall grid emissions. Carbon accounting for ES systems has attracted considerable recent attention from the industry [34, 35]. References [29, 28, 36] propose different carbon footprint models for ES systems, while they neglect the carbon emission leakage associated with ES energy loss (see Remark 4), and the carbon accounting mechanisms for ES system owners remain unclear.
@@ II-B Carbon Emission Flow Model for Lossy Power Networks
Equation (4) is referred to as the carbon flow model or carbon flow equation, and it can be equivalently reformulated in the matrix form (5). The detailed derivation of (5a) is provided in Appendix A.
$(\bm{P}_{\mathrm{N}}-\bm{P}_{\mathrm{B}})\bm{w}_{\mathrm{N}}=\bm{r}_{\mathrm{G}}$ (5a)
$\qquad\Longrightarrow\ \bm{w}_{\mathrm{N}}=(\bm{P}_{\mathrm{N}}-\bm{P}_{\mathrm{B}})^{-1}\bm{r}_{\mathrm{G}},$ (5b)
Here, $\bm{w}_{\mathrm{N}}\!:=\!(w_{i})_{i\in\mathcal{N}}\!\in\!\mathbb{R}^{N}$ and $\bm{r}_{\mathrm{G}}\!:=\!(\sum_{g\in\mathcal{G}_{i}}\!\!w_{i,g}^{\mathrm{G}}P_{i,g}^{\mathrm{G}})_{i\in\mathcal{N}}\!\in\!\mathbb{R}^{N}$ denote the column vectors that collect the nodal carbon intensities and nodal generation carbon flows, respectively. $\bm{P}_{\mathrm{N}}\!:=\!\text{diag}(P_{i}^{\mathrm{in}})\!\in\!\mathbb{R}^{N\times N}$ is the diagonal matrix whose $i$-th diagonal entry is the nodal active power inflow $P_{i}^{\mathrm{in}}$ to node $i$. $\bm{P}_{\mathrm{B}}\!\in\!\mathbb{R}^{N\times N}$ is the branch power inflow matrix that is built by letting $\bm{P}_{\mathrm{B}}[i,k]\!=\!P_{ki}^{i}$ and $\bm{P}_{\mathrm{B}}[k,i]\!=\!0$ if node $k$ sends power flow $P_{ki}^{i}$ to node $i$. Note that $P_{ki}^{i}$ and $P_{ki}^{k}$ are different due to the line power loss. Given power flow profiles, the linear equations (5a) can be solved to obtain $\bm{w}_{\mathrm{N}}$, e.g., through matrix factorization and backward and forward substitution, rather than directly computing the matrix inverse in (5b).
@@ II-C Feasibility and Solution Uniqueness of Carbon Flow Model
According to [39, Theorem], the matrix $\bm{P}_{\mathrm{C}}$ is nonsingular and thus invertible. Since $\bm{P}_{\mathrm{C}}$ is diagonally dominant and all its off-diagonal elements $\{\bm{P}_{\mathrm{C}}[i,j]\}_{i\neq j}$ are non-positive, $\bm{P}_{\mathrm{C}}$ is an M-matrix [40] according to [39, Corollary 4]. ∎
@@
The condition in Theorem 1 can be interpreted as that for any node $i$ with no generation power injection, one can find a power flow path $i\leftarrow i_{1}\leftarrow i_{2}\leftarrow\cdots\leftarrow i_{r}\leftarrow j$ in the power network that traces upstream to a node $j$ that has generation power injection. This condition generally holds for practical connected power networks. This condition also implies that every node has non-zero power flux, i.e., $\bm{P}_{\mathrm{C}}[i,i]>0$ for all $i\in[N]$. However, non-zero nodal power flux can not guarantee that $\bm{P}_{\mathrm{C}}$ is invertible. For example, consider a simple 3-node circular network case shown in Figure 3. In this case, every node has 1 unit of power flowing through it, but the matrix $\bm{P}_{\mathrm{C}}$ is singular. And this case violates the condition in Theorem 1. In addition, $\bm{P}_{\mathrm{C}}$ possesses all the properties of being an M-matrix. For example, all the elements of $\bm{P}_{\mathrm{C}}^{-1}$ are non-negative and every eigenvalue of $\bm{P}_{\mathrm{C}}$ has a positive real part [40, Theorem 1.1].
@@ III-A1 Objective Function
where (8a) denotes the total generation cost in a quadratic form with the parameters $c_{i,g}^{2},c_{i,g}^{1},c_{i,g}^{0}$, and (8b) is the penalty on generation-side emissions with the cost coefficient $c^{\mathrm{emi}}$.
@@ III-A2 Power Flow Equations and Constraints
The power flow constraints (7c) generally involve the line thermal capacity constraints (10a) and voltage limits (10b).
@@
where $\bar{S}_{ij}$ is the apparent power capacity of line $ij$. $\bar{V}_{i},\underline{V}_{i}$ are the upper and lower limits of voltage magnitude at node $i$.
@@ III-A3 Carbon Flow Equations and Constraints
Essentially, the C-OPF model (7) is a carbon-aware generalization of the OPF model, and it reduces to an OPF model if the carbon-related objective function $f_{\mathrm{carbon}}$, carbon flow equations (7d) and constraints (7e) are removed or inactive. It also implies that existing OPF techniques, such as linearization, convexification, decomposition, stochastic modeling, etc., can still be applied to the power flow components in C-OPF (7). Moreover, the C-OPF model (7) can be directly extended to the multi-period settings and involve time-coupled constraints, such as generator ramping limits and the state-of-charge constraints of ES systems. As an illustrative example, a multi-period C-OPF-based economic dispatch model (27) is established in Section V-A, which includes these time-coupled constraints and employs a full AC power flow model.
In the C-OPF model (7), the carbon flow method is used for demand-side carbon accounting to align the grid’s carbon footprint quantification with physical power system operation and actual power flows. Nevertheless, the C-OPF method is flexible and can adapt to other carbon accounting approaches by replacing the carbon flow equation (7d) with other valid carbon accounting mathematical models.
@@ IV Carbon Footprint Model for Energy Storage
Energy storage (ES) systems play a critical role in decarbonizing power grids, as their operations can be optimized to curtail overall system emissions, e.g., charging when the grid is clean and discharging when the grid is under high emissions. Consider an ES system connected to node $i\!\in\!\mathcal{N}$. For time $t\in\mathcal{T}:=\{1,2,\cdots,T\}$ with the time interval $\delta_{t}$, the dynamical ES power model [45] is formulated as (20):
@@ V-C1 Nodal Carbon Intensity
Figure 6 illustrates the nodal carbon intensities for all load nodes under the C-OPF-based and OPF-based ED schemes, respectively. It shows that the C-OPF-based ED model can generate effective power dispatch schemes that keep the nodal carbon intensities of load nodes below the cap of $\bar{w}_{i,t}\!=\!1.2$ lbs/kWh. In contrast, the OPF-based ED scheme frequently exceeds the nodal carbon intensity cap. Moreover, the grid’s nodal carbon intensities exhibit significant temporal variation and spatial diversity. From Figure 6, it is observed that the nodal carbon intensities at 13:00 are generally lower than those at 21:00, due to the higher penetration of renewable generation at 13:00, as illustrated in Figure 8. Figure 7 visualizes the grid’s nodal carbon intensities at 13:00, with darker blue colors indicating higher carbon intensity at each node. It demonstrates that the nodal carbon intensities calculated using the carbon flow method can reflect the proximity to different fuel types of generation and align with the physical power flow. In comparison, the grid average emission factor at 13:00 is calculated to be 0.5 lbs/kWh, which only gives an indication of the overall grid average emission state and falls short of providing detailed insight into local emissions at different locations.
@@ V-F Computational Efficiency
Additionally, we implement the DC power flow model (11) instead of the full AC power flow model in the C-OPF-based and OPF-based ED models for simulation comparison. In this case, both ED models are significantly simplified due to omitting voltage magnitude and reactive power, neglecting power line losses, and replacing nonlinear power flow equations with linear DC flow equations. Under the DC power flow model, it takes 16.4 seconds on average to solve the C-OPF-based ED model and 1.2 seconds to solve the OPF-based ED model.
Due to the dual power flow reformulation and the addition of carbon flow equations and constraints, the C-OPF model generally has a larger problem size and requires more solution time than the traditional OPF model. In practice, several methods can improve the solution efficiency of the C-OPF model. For instance, by inspecting all branches and identifying those whose power flow directions can be predetermined based on the network configuration, the dual power flow reformulation can be avoided for these branches. Additionally, the solutions of the OPF model can be employed to warm-start the C-OPF model. A key future research direction is to develop efficient linearization and convexification approaches for the nonconvex carbon flow equations to fundamentally enhance the solution efficiency of the C-OPF model.